# The Great Data Standoff: Researchers vs. Platforms Under the Digital Services Act

### v1

@@ 3.2. Platform Manipulation and Election Interference: Related Studies and Data Needs
Specifically, Shao et al. (Shao et al., 2018) analyse social bots and their role in sharing low-credibility content, finding that during the 2016 US elections, social bots shared a disproportionate amount of low-credibility content. To conduct their study, they used the following data: 1) Content Data (tweets), 2) User data (Twitter account information), 3) Moderation Data (news sources with low credibility information), and 4) Engagement Data (retweet information). Bessi and Ferrara (Bessi and Ferrara, 2016) investigate how social bots affect political discussion during the 2016 US elections, finding that the presence of bots can negatively affect political discussion on Twitter. To conduct their research, they relied on the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Temporal Data (tweets over time), and 4) Geospatial Data (location of the tweets). The same kind of data was also applied by Ferrara (Ferrara, 2017) to study social bots during the 2017 French election. Ferrara et al. (Ferrara et al., 2020) focus on the 2020 US elections and show that social bots online exacerbate the problem of political echo chambers. For their work, they used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Temporal Data (tweets over time), 4) Moderation Data (tweets shared by state-sponsored actors), and 5) Engagement Data (retweet information).
@@
Another well-known implementation of such campaigns uses state-sponsored trolls, which are humans who operate a set of fake accounts, are sponsored by their governments, and aim to push specific narratives on social media to push their government’s agenda. Examples include Russian trolls active on social media platforms during the 2016 US elections (Badawy et al., 2018). Previous work on coordinated campaigns using state-sponsored trolls focused on analysing their characteristics and the strategies employed (Zannettou et al., 2019a), detecting actors involved in such campaigns during elections (Luceri et al., 2020), and understanding their impact and influence on the Web and our society (Golovchenko et al., 2020; Badawy et al., 2018). Specifically, Badawy et al. (Badawy et al., 2018) investigate the Russian interference campaign on Twitter during the 2016 US presidential election. They analysed a large dataset of tweets to demystify the strategies used by the Internet Research Agency (IRA), a Russian organization known for its online propaganda, and analysed the Twitter users that engaged with tweets shared by IRA. The data used for this study include: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Network Data (retweet network), 4) Moderation Data (Accounts identified as IRA accounts), 5) Temporal Data (IRA tweets over time), 6) Engagement Data (people who engaged with IRA tweets), and 7) Geospatial Data (Location of tweet posting). Zannettou et al. (Zannettou et al., 2019a) analyse the behaviour and strategies employed by IRA trolls on Twitter during the 2016 US elections and assess the influence that they had on other platforms like 4chan and Reddit. This study leveraged the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Moderation Data (Accounts identified as IRA accounts), 4) Temporal Data (IRA tweets over time), and 5) Geospatial Data (Location of tweet posting). Similarly, Golovchenko et al. (Golovchenko et al., 2020) investigated IRA’s tweets during the 2016 US elections, focusing on analysing the YouTube videos included in IRA tweets. The data used for this study include: 1) Content data (tweets), 2) User data (Twitter account information), and 3) Moderation data (Accounts identified as IRA accounts). Other research efforts focused on detecting actors involved in such campaigns. An example is the work from Luceri et al. (Luceri et al., 2020), which focused on detecting state-sponsored trolls that were involved in the 2016 US elections using an inverse reinforcement learning approach. To achieve this, they used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), and 3) Moderation Data (Accounts identified as IRA accounts).
@@
Another method of recruiting user accounts for promoting the messages of a coordinated information campaign is employing social-media influencers. This strategy has been observed soon after Russia’s invasion of Ukraine (“Russian lives matter” campaign (Richards, 2022)) and has also been noted in the Romanian 2024 presidential elections (Popoviciu, 2024). We discuss more on this strategy in the context of hidden advertising in Section 3.3.
@@
Identifying Algorithmic Manipulation. Algorithmic manipulation refers to the intentional exploitation of platform algorithms and systems to amplify/suppress specific narratives or distort user engagement (Conti et al., 2024). Platform algorithms include recommendation systems (systems responsible for recommending content to users), algorithms for content search and ranking, and algorithms that identify trending topics based on user activity (Elmas et al., 2021). Nowadays, social media platforms increasingly rely on AI algorithms to enhance user experience, including recommendation systems to generate personalised feeds, recommend users to follow, etc. Given the increasing reliance of social media platforms on AI-based algorithms for recommending content to users, there are several concerns with respect to how such systems can be exploited by malicious actors that aim to manipulate the public and promote/suppress specific narratives online.
@@
Motivated by this, previous research focused on understanding these phenomena. For instance, Conti et al. (Conti et al., 2024) conducted a quantitative study on suspicious changes in content visibility on Twitter, likely due to algorithmic interventions. To conduct their research, the authors used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Algorithmic Data (content visibility as determined by the algorithm ), 4) Temporal Data (tweets over time), 5) Engagement Data (retweet information), and 6) Moderation Data (content factuality based on the sources). Another example of algorithmic manipulation was shown by Elmas et al. (Elmas et al., 2021), who demonstrated how the Twitter trending algorithm can be manipulated to promote specific political content in Turkey during the 2019 Istanbul elections. To demonstrate this kind of algorithmic manipulation, they used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Algorithmic Data (Trending topics), and 4) Temporal Data (algorithmic manipulation attacks over time).
@@
  • Content Data: Text, images, videos, sound, transcriptions, hashtags, user mentions (Allcott and Gentzkow, 2017; Grinberg et al., 2019; Zannettou, 2021). Example use case: Analyse and detect posts that share false information during elections.
@@
  • User Data: Account metadata, such as the account creation date, follower count, user demographics, bios, status, etc. Example use case: Study the personas of accounts pertaining to state-sponsored actors (Badawy et al., 2018; Zannettou et al., 2019a; Luceri et al., 2020).
@@
  • Engagement Data: Metrics such as likes, shares, comments, and views (Allcott and Gentzkow, 2017; Bovet and Makse, 2019; Zannettou, 2021). Example use case: Analysing user engagement with false information related to elections.
@@
  • Moderation Data: Posts or content flagged, removed, or labelled by the platform or external sources (e.g., fact-checkers) (Badawy et al., 2018; Luceri et al., 2020; Zannettou et al., 2019a; Ferrara et al., 2020; Shao et al., 2018). Also, moderation data on the user level, e.g. whether a user account is suspected to be a bot or controlled by a state-sponsored actor. Example use case: Studying the effectiveness of platform moderation interventions for posts sharing false information.
@@
  • Temporal Data: Time-based activity logs, such as the engagement that a misinformation record receives over time (Grinberg et al., 2019; Zannettou, 2021; Conti et al., 2024). Example use case: Analysing a timeline of the dissemination of a false narrative before an election.
@@
  • Network Data: Data about the social graph, such as user connections and interactions. Example use case: Analysing influential nodes in disseminating false information (Grinberg et al., 2019; Badawy et al., 2018).
@@
  • Algorithmic Data: Data on algorithmic outputs and decisions (e.g., videos recommended to end-users), as well as the importance of various features in algorithmic outputs (Elmas et al., 2021; Conti et al., 2024). Example use case: Investigating changes in content visibility due to algorithmic manipulation.
@@
  • Geospatial Data: Anonymised and aggregate location data (Badawy et al., 2018; Zannettou et al., 2019a; Bessi and Ferrara, 2016; Ferrara, 2017). Example use case: Identifying and analysing coordinated campaigns originating from specific regions.
3.3. Political Influencer Marketing as Hidden Advertising
@@
The popularity of influencer marketing has been growing steadily from US\$35.1bn in 2024 to US\$52bn expected in 2028 (Statista, [n. d.]). Its advantages, particularly the opportunity to embed advertising with organic content made by creators who establish parasocial relationships with their followers, make it an incredibly compelling marketing strategy – even for political advertising (Banjo, [n. d.]). Candidates and political entities find it important to partake in the creator economy, and creators themselves are pressured to make more political content and engage with social justice questions (Siu, [n. d.]).
@@
Measuring the Prevalence of Undisclosed Ads. Undisclosed advertisements have been widely acknowledged as a major regulatory concern by various legislative bodies, including the European Commission (EC) drafting the DSA (Duivenvoorde and Goanta, 2023). Enforcement of the DSA in the realm of political advertising is of great significance since it could safeguard the free expression of diverse political ideologies and help maintain democratic integrity. In the context of the Romanian elections, political ideologies were disseminated through seemingly credible media channels, including social media influencers. When political advertising is not disclosed, audiences are less likely to recognise the content as advertisements (or other paid arrangements), leaving them vulnerable to manipulation (Eisend et al., 2020; Liu et al., 2024). This lack of transparency allows certain political candidates to exploit misinformation and covert persuasion tactics, potentially swaying voter behaviour and gaining an unfair electoral advantage. Such practices not only undermine voter autonomy but also jeopardise the legitimacy of democratic processes. This issue aligns with broader concerns regarding systemic risks posed by hidden advertising as illegal content. Measuring the prevalence of hidden political advertisements is, therefore, a critical step in establishing whether they are a factor in the understanding of election-related systemic risks. By assessing how frequently and in what forms undisclosed political ads appear, regulators can gain valuable insights into the scope of the problem. This, in turn, can inform the development of targeted interventions to ensure elections remain fair, transparent, and free from manipulation, thereby upholding democratic values and fostering trust in political elections.
@@
Measuring the prevalence of undisclosed ads first requires identifying which posts contain sponsored content. Existing sponsored content detection research has primarily focused on commercial content, framing the problem as a machine learning classification task. Many models adopt a semi-supervised approach, using disclosure hashtags or keywords (e.g., #ad, #sponsored) as weak labels to distinguish between sponsored and organic posts (Kim et al., 2021; Bertaglia et al., 2023). Studies rely primarily on Content Data (such as post captions, engagement metrics, and associated metadata) to train models (Zarei et al., 2020; Bertaglia et al., 2023, 2025). Kim et al. (2021) incorporate Network Data to model the relationship between influencers, brands, and (sponsored) posts and add more context to the machine learning models. Such approaches require additional User Data, including account metadata (e.g., follower counts, account type). Other studies combine image and text data to improve model performance (Sánchez Villegas et al., 2023; Villegas et al., 2021). In this context, such multimodal models would benefit from additional Content Data, including attributes inferred by the platforms (e.g., detected objects in videos, audio transcriptions, and visual markers of sponsorship).
@@
Measuring the Engagement of Organic Content, Disclosed and Undisclosed Ads. A key reason behind the non-disclosure of advertisements is the widespread belief that revealing sponsorships harms audience engagement. Influencers, who simultaneously act as content creators and advertising platforms, grapple with the tension between protecting their long-term growth and maximizing short-term profits. The assumption that disclosure reduces engagement incentivises some influencers to conceal sponsorships. Understanding the actual impact of advertisements and their (non)disclosure on engagement is therefore crucial for assessing the validity of this belief. Empirical studies have analysed the effects of disclosures and regulations on engagement, relying on Engagement Data to compare audience reactions to disclosed versus undisclosed ads, as well as before and after specific regulations (Ershov and Mitchell, 2020; Waltenrath, 2021; Bertaglia et al., 2025). However, effectively measuring these differences requires a longitudinal perspective; therefore, Temporal Data, especially in the form of engagement metrics over time, is essential for understanding the persistence and impact of undisclosed advertising. Research can be designed to compare engagement levels across organic content and disclosed and undisclosed advertisements. If empirical findings refute the assumption that disclosure negatively affects engagement, this could pave the way for educational initiatives aimed at influencers, emphasising that transparency does not undermine engagement. Such efforts could, in turn, improve compliance rates with disclosure regulations.
@@
Assessing How Platforms Promote or Moderate Sponsored Content. Another reason for the non-disclosure of advertisements stems from a belief among influencers that platforms deliberately restrict the visibility of advertising content (Musiyiwa and Jacobson, 2023; Savolainen, 2022; Bishop, 2019). This perception is linked to the inherent tension platforms face between increasing user traffic and maximising financial gains. While displaying more ads aligns with the interests of advertisers, it risks alienating users and discouraging them from engaging with the platform. Consequently, platforms are compelled to strike a balance between promoting and moderating advertisements to sustain their ecosystem. Due to this concern, an ideal choice for the influencers would be to hide the advertisement so that they can bypass the moderation process. To address this issue, it is crucial to understand how platforms’ content promotion and moderation mechanisms operate. Research could explore the factors influencing the visibility of advertising content, such as algorithmic preferences, user behaviour data and the balance between paid and organic content. By examining the biases inherent in these systems, such research could identify patterns that favour or limit certain types of content, providing valuable insights into how platforms manage the promotion and visibility of advertisements. In practice, investigating whether and how platforms moderate or amplify advertising content requires access to several key data types. In addition to the data points identified for the previous tasks, Algorithmic Data, such as content recommended to users and features related to their ranking, would be essential for determining how sponsored content performs in the recommender system within the platform.
@@
  • Content Data: All media content (text, images, videos, and sound), metadata (e.g., captions, timestamps, and hashtags), and inferred attributes (e.g., object detection and visual sponsorship markers) from influencer posts. These data points are fundamental for identifying advertising content (Zarei et al., 2020; Bertaglia et al., 2023; Villegas et al., 2021; Sánchez Villegas et al., 2023). Example use case: Training machine learning models to detect undisclosed political ads.
  • User Data: Metadata related to influencer accounts, including follower counts, account creation date, account type, and inferred demographics. Studying influencer engaging in political campaigns requires access to lists of content creators and their attributes (Kim et al., 2021; Bertaglia et al., 2025). Example use case: Mapping Romanian influencers involved in political campaigns by analysing account demographics and activity.
  • Engagement Data: Metrics such as likes, shares, comments, and views across influencer posts (Ershov and Mitchell, 2020; Waltenrath, 2021). Example use case: Analysing the impact of disclosures and regulation in engagement with political posts.
  • Moderation Data: Information on platform-inferred labels, content flagging, removals, and downranking of sponsored and political content. This data would help determine whether automated moderation disproportionately affects disclosed versus undisclosed ads (Musiyiwa and Jacobson, 2023; Savolainen, 2022). Example use case: Assessing whether platform moderation algorithms label disclosed political ads differently from undisclosed ones.
  • Temporal Data: Engagement metrics over time (Ershov and Mitchell, 2020; Bertaglia et al., 2025). Example use case: Identify potential coordinated engagement with political content.
  • Network Data: Interaction data capturing connections between influencers, brands, political entities, and audience members. Mapping these relationships helps detect coordinated campaigns and assessing the diffusion of undisclosed ads (Kim et al., 2021). Example use case: Detecting coordinated influence operations by analysing how content spreads across influencer networks.
  • Algorithmic Data: Includes recommendation system outputs, ranking metrics, and visibility logs for influencer content. Access to this data is essential to assess whether platform recommendation systems amplify or suppress disclosed advertisements compared to undisclosed content (Savolainen, 2022). Example use case: Analysing whether recommendation algorithms rank disclosed political ads differently from undisclosed ones.
4. DATA AVAILABILITY AS A PLATFORM GOVERNANCE NARRATIVE

### v2

@@ A Brief Overview of an Election Case Study
Specifically, Shao et al. (shao2018spread) analyse social bots and their role in sharing low-credibility content, finding that during the 2016 US elections, social bots shared a disproportionate amount of low-credibility content. To conduct their study, they used the following data: 1) Content Data (tweets), 2) User data (Twitter account information), 3) Moderation Data (news sources with low credibility information), and 4) Engagement Data (retweet information). Bessi and Ferrara (bessi2016social) investigate how social bots affect political discussion during the 2016 US elections, finding that the presence of bots can negatively affect political discussion on Twitter. To conduct their research, they relied on the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Temporal Data (tweets over time), and 4) Geospatial Data (location of the tweets). The same kind of data was also applied by Ferrara (ferrara2017disinformation) to study social bots during the 2017 French election. Ferrara et al. (ferrara2020characterizing) focus on the 2020 US elections and show that social bots online exacerbate the problem of political echo chambers. For their work, they used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Temporal Data (tweets over time), 4) Moderation Data (tweets shared by state-sponsored actors), and 5) Engagement Data (retweet information).
@@
Another well-known implementation of such campaigns uses state-sponsored trolls, which are humans who operate a set of fake accounts, are sponsored by their governments, and aim to push specific narratives on social media to push their government’s agenda. Examples include Russian trolls active on social media platforms during the 2016 US elections (badawy2018analyzing). Previous work on coordinated campaigns using state-sponsored trolls focused on analysing their characteristics and the strategies employed (zannettou2019disinformation), detecting actors involved in such campaigns during elections (luceri2020detecting), and understanding their impact and influence on the Web and our society (golovchenko2020cross; badawy2018analyzing). Specifically, Badawy et al. (badawy2018analyzing) investigate the Russian interference campaign on Twitter during the 2016 US presidential election. They analysed a large dataset of tweets to demystify the strategies used by the Internet Research Agency (IRA), a Russian organization known for its online propaganda, and analysed the Twitter users that engaged with tweets shared by IRA. The data used for this study include: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Network Data (retweet network), 4) Moderation Data (Accounts identified as IRA accounts), 5) Temporal Data (IRA tweets over time), 6) Engagement Data (people who engaged with IRA tweets), and 7) Geospatial Data (Location of tweet posting). Zannettou et al. (zannettou2019disinformation) analyse the behaviour and strategies employed by IRA trolls on Twitter during the 2016 US elections and assess the influence that they had on other platforms like 4chan and Reddit. This study leveraged the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Moderation Data (Accounts identified as IRA accounts), 4) Temporal Data (IRA tweets over time), and 5) Geospatial Data (Location of tweet posting). Similarly, Golovchenko et al. (golovchenko2020cross) investigated IRA’s tweets during the 2016 US elections, focusing on analysing the YouTube videos included in IRA tweets. The data used for this study include: 1) Content data (tweets), 2) User data (Twitter account information), and 3) Moderation data (Accounts identified as IRA accounts). Other research efforts focused on detecting actors involved in such campaigns. An example is the work from Luceri et al. (luceri2020detecting), which focused on detecting state-sponsored trolls that were involved in the 2016 US elections using an inverse reinforcement learning approach. To achieve this, they used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), and 3) Moderation Data (Accounts identified as IRA accounts).
@@
Another method of recruiting user accounts for promoting the messages of a coordinated information campaign is employing social-media influencers. This strategy has been observed soon after Russia’s invasion of Ukraine (“Russian lives matter” campaign (richards_pro-russia_2022)) and has also been noted in the Romanian 2024 presidential elections (popoviciu_tiktok_2024). We discuss more on this strategy in the context of hidden advertising in Section Political Influencer Marketing as Hidden Advertising.
@@
Identifying Algorithmic Manipulation. Algorithmic manipulation refers to the intentional exploitation of platform algorithms and systems to amplify/suppress specific narratives or distort user engagement (conti2024revealing). Platform algorithms include recommendation systems (systems responsible for recommending content to users), algorithms for content search and ranking, and algorithms that identify trending topics based on user activity (elmas2021ephemeral). Nowadays, social media platforms increasingly rely on AI algorithms to enhance user experience, including recommendation systems to generate personalised feeds, recommend users to follow, etc. Given the increasing reliance of social media platforms on AI-based algorithms for recommending content to users, there are several concerns with respect to how such systems can be exploited by malicious actors that aim to manipulate the public and promote/suppress specific narratives online.
@@
Motivated by this, previous research focused on understanding these phenomena. For instance, Conti et al. (conti2024revealing) conducted a quantitative study on suspicious changes in content visibility on Twitter, likely due to algorithmic interventions. To conduct their research, the authors used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Algorithmic Data (content visibility as determined by the algorithm ), 4) Temporal Data (tweets over time), 5) Engagement Data (retweet information), and 6) Moderation Data (content factuality based on the sources). Another example of algorithmic manipulation was shown by Elmas et al. (elmas2021ephemeral), who demonstrated how the Twitter trending algorithm can be manipulated to promote specific political content in Turkey during the 2019 Istanbul elections. To demonstrate this kind of algorithmic manipulation, they used the following data: 1) Content Data (tweets), 2) User Data (Twitter account information), 3) Algorithmic Data (Trending topics), and 4) Temporal Data (algorithmic manipulation attacks over time).
@@
  • Content Data: Text, images, videos, sound, transcriptions, hashtags, user mentions (allcott2017social; grinberg2019fake; zannettou2021won). Example use case: Analyse and detect posts that share false information during elections.
@@
  • User Data: Account metadata, such as the account creation date, follower count, user demographics, bios, status, etc. Example use case: Study the personas of accounts pertaining to state-sponsored actors (badawy2018analyzing; zannettou2019disinformation; luceri2020detecting).
@@
  • Engagement Data: Metrics such as likes, shares, comments, and views (allcott2017social; bovet2019influence; zannettou2021won). Example use case: Analysing user engagement with false information related to elections.
@@
  • Moderation Data: Posts or content flagged, removed, or labelled by the platform or external sources (e.g., fact-checkers) (badawy2018analyzing; luceri2020detecting; zannettou2019disinformation; ferrara2020characterizing; shao2018spread). Also, moderation data on the user level, e.g. whether a user account is suspected to be a bot or controlled by a state-sponsored actor. Example use case: Studying the effectiveness of platform moderation interventions for posts sharing false information.
@@
  • Temporal Data: Time-based activity logs, such as the engagement that a misinformation record receives over time (grinberg2019fake; zannettou2021won; conti2024revealing). Example use case: Analysing a timeline of the dissemination of a false narrative before an election.
@@
  • Network Data: Data about the social graph, such as user connections and interactions. Example use case: Analysing influential nodes in disseminating false information (grinberg2019fake; badawy2018analyzing).
@@
  • Algorithmic Data: Data on algorithmic outputs and decisions (e.g., videos recommended to end-users), as well as the importance of various features in algorithmic outputs (elmas2021ephemeral; conti2024revealing). Example use case: Investigating changes in content visibility due to algorithmic manipulation.
@@
  • Geospatial Data: Anonymised and aggregate location data (badawy2018analyzing; zannettou2019disinformation; bessi2016social; ferrara2017disinformation). Example use case: Identifying and analysing coordinated campaigns originating from specific regions.
Political Influencer Marketing as Hidden Advertising
@@
The popularity of influencer marketing has been growing steadily from US\$35.1bn in 2024 to US\$52bn expected in 2028 (statista_influencer_2024). Its advantages, particularly the opportunity to embed advertising with organic content made by creators who establish parasocial relationships with their followers, make it an incredibly compelling marketing strategy – even for political advertising. Candidates and political entities find it important to partake in the creator economy, and creators themselves are pressured to make more political content and engage with social justice questions (siu_how_2024).
@@
Measuring the Prevalence of Undisclosed Ads. Undisclosed advertisements have been widely acknowledged as a major regulatory concern by various legislative bodies, including the European Commission (EC) drafting the DSA (duivenvoorde2023regulation). Enforcement of the DSA in the realm of political advertising is of great significance since it could safeguard the free expression of diverse political ideologies and help maintain democratic integrity. In the context of the Romanian elections, political ideologies were disseminated through seemingly credible media channels, including social media influencers. When political advertising is not disclosed, audiences are less likely to recognise the content as advertisements (or other paid arrangements), leaving them vulnerable to manipulation (liu2024effects). This lack of transparency allows certain political candidates to exploit misinformation and covert persuasion tactics, potentially swaying voter behaviour and gaining an unfair electoral advantage. Such practices not only undermine voter autonomy but also jeopardise the legitimacy of democratic processes. This issue aligns with broader concerns regarding systemic risks posed by hidden advertising as illegal content. Measuring the prevalence of hidden political advertisements is, therefore, a critical step in establishing whether they are a factor in the understanding of election-related systemic risks. By assessing how frequently and in what forms undisclosed political ads appear, regulators can gain valuable insights into the scope of the problem. This, in turn, can inform the development of targeted interventions to ensure elections remain fair, transparent, and free from manipulation, thereby upholding democratic values and fostering trust in political elections.
@@
Measuring the prevalence of undisclosed ads first requires identifying which posts contain sponsored content. Existing sponsored content detection research has primarily focused on commercial content, framing the problem as a machine learning classification task. Many models adopt a semi-supervised approach, using disclosure hashtags or keywords (e.g., #ad, #sponsored) as weak labels to distinguish between sponsored and organic posts (kimDiscoveringUndisclosedPaid2021b; bertaglia2023closing). Studies rely primarily on Content Data (such as post captions, engagement metrics, and associated metadata) to train models (zareiCharacterisingDetectingSponsored2020; bertaglia2023closing; bertaglia2025influencer). kimDiscoveringUndisclosedPaid2021b incorporate Network Data to model the relationship between influencers, brands, and (sponsored) posts and add more context to the machine learning models. Such approaches require additional User Data, including account metadata (e.g., follower counts, account type). Other studies combine image and text data to improve model performance (sanchez-villegas-etal-2023-multimodal; villegas2021analyzing). In this context, such multimodal models would benefit from additional Content Data, including attributes inferred by the platforms (e.g., detected objects in videos, audio transcriptions, and visual markers of sponsorship).
@@
Measuring the Engagement of Organic Content, Disclosed and Undisclosed Ads. A key reason behind the non-disclosure of advertisements is the widespread belief that revealing sponsorships harms audience engagement. Influencers, who simultaneously act as content creators and advertising platforms, grapple with the tension between protecting their long-term growth and maximising short-term profits. The assumption that disclosure reduces engagement incentivises some influencers to conceal sponsorships. Understanding the actual impact of advertisements and their (non)disclosure on engagement is therefore crucial for assessing the validity of this belief. Empirical studies have analysed the effects of disclosures and regulations on engagement, relying on Engagement Data to compare audience reactions to disclosed versus undisclosed ads, as well as before and after specific regulations (ershovEffectsInfluencerAdvertising2020; bertaglia2025influencer). However, effectively measuring these differences requires a longitudinal perspective; therefore, Temporal Data, especially in the form of engagement metrics over time, is essential for understanding the persistence and impact of undisclosed advertising. Research can be designed to compare engagement levels across organic content and disclosed and undisclosed advertisements. If empirical findings refute the assumption that disclosure negatively affects engagement, this could pave the way for educational initiatives aimed at influencers, emphasising that transparency does not undermine engagement. Such efforts could, in turn, improve compliance rates with disclosure regulations.
@@
Assessing How Platforms Promote or Moderate Sponsored Content. Another reason for the non-disclosure of advertisements stems from a belief among influencers that platforms deliberately restrict the visibility of advertising content (musiyiwa2023sponsorship; savolainen2022shadow; bishop2019managing). This perception is linked to the inherent tension platforms face between increasing user traffic and maximising financial gains. While displaying more ads aligns with the interests of advertisers, it risks alienating users and discouraging them from engaging with the platform. Consequently, platforms are compelled to strike a balance between promoting and moderating advertisements to sustain their ecosystem. Due to this concern, an ideal choice for the influencers would be to hide the advertisement so that they can bypass the moderation process. To address this issue, it is crucial to understand how platforms’ content promotion and moderation mechanisms operate. Research could explore the factors influencing the visibility of advertising content, such as algorithmic preferences, user behaviour data and the balance between paid and organic content. By examining the biases inherent in these systems, such research could identify patterns that favour or limit certain types of content, providing valuable insights into how platforms manage the promotion and visibility of advertisements. In practice, investigating whether and how platforms moderate or amplify advertising content requires access to several key data types. In addition to the data points identified for the previous tasks, Algorithmic Data, such as content recommended to users and features related to their ranking, would be essential for determining how sponsored content performs in the recommender system within the platform.
@@
  • Content Data: All media content (text, images, videos, and sound), metadata (e.g., captions, timestamps, and hashtags), and inferred attributes (e.g., object detection and visual sponsorship markers) from influencer posts. These data points are fundamental for identifying advertising content (zareiCharacterisingDetectingSponsored2020; bertaglia2023closing; villegas2021analyzing; sanchez-villegas-etal-2023-multimodal). Example use case: Training machine learning models to detect undisclosed political ads.
  • User Data: Metadata related to influencer accounts, including follower counts, account creation date, account type, and inferred demographics. Studying influencer engaging in political campaigns requires access to lists of content creators and their attributes (kimDiscoveringUndisclosedPaid2021b; bertaglia2025influencer). Example use case: Mapping Romanian influencers involved in political campaigns by analysing account demographics and activity.
  • Engagement Data: Metrics such as likes, shares, comments, and views across influencer posts (ershovEffectsInfluencerAdvertising2020). Example use case: Analysing the impact of disclosures and regulation in engagement with political posts.
  • Moderation Data: Information on platform-inferred labels, content flagging, removals, and downranking of sponsored and political content. This data would help determine whether automated moderation disproportionately affects disclosed versus undisclosed ads (musiyiwa2023sponsorship; savolainen2022shadow). Example use case: Assessing whether platform moderation algorithms label disclosed political ads differently from undisclosed ones.
  • Temporal Data: Engagement metrics over time (ershovEffectsInfluencerAdvertising2020; bertaglia2025influencer). Example use case: Identify potential coordinated engagement with political content.
  • Network Data: Interaction data capturing connections between influencers, brands, political entities, and audience members. Mapping these relationships helps detect coordinated campaigns and assessing the diffusion of undisclosed ads (kimDiscoveringUndisclosedPaid2021b). Example use case: Detecting coordinated influence operations by analysing how content spreads across influencer networks.
  • Algorithmic Data: Includes recommendation system outputs, ranking metrics, and visibility logs for influencer content. Access to this data is essential to assess whether platform recommendation systems amplify or suppress disclosed advertisements compared to undisclosed content (savolainen2022shadow). Example use case: Analysing whether recommendation algorithms rank disclosed political ads differently from undisclosed ones.
Data Availability as a Platform Governance Narrative